# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5.3.4 Benchmark models
Tab. LABEL:t:benchmark_models_finance displays the benchmark models utilized in the ”Finance” group by a minimum of two of the selected papers.
Table 11: List of benchmark models in the ”Finance” group divided per category.
@@
Mathematical and statistical methods Mean reversion (MR) [art_63], [art_65]
Classical machine learning methods Convolutional neural network (CNN) [art_59], [art_64]
Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65]
Feed-forward neural network (FNN) [art_59], [art_68]
Gated recurrent unit (GRU) [art_60], [art_68]
Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68]
Support vector machine (SVM) [art_64], [art_68]
GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68]
@@
Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65]
Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65]
It is evident that there is no consensus regarding the most appropriate benchmark models, especially in the context of GNNs. The most common benchmark is the LSTM, which is considered an effective approach for processing long term information. The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al. in 2019 [HATS]. The HATS model has been proposed for predicting stock movements using relational data. It aggregates information from different types of relationships among data and incorporates this information into each stock representation. The source code for the HATS model is available at https://github.com/dmis-lab/hats.
@@ 5.3.5 Results
The accuracy of the proposed classification models is primarily evaluated in terms of accuracy, precision and F1 score. The only two papers with comparable results are Refs. [art_63] and [art_65], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock). These two papers compare the results obtained by using 5, 10 and 20 trading days as records. For sake of simplicity, a comparison of the accuracy of the models in the two papers with 20 days only is provided in Tab. LABEL:t:risultati_a-share.
@@
The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [art_63] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [art_65], depending on the metrics considered.
@@ 5.4 Health
The fourth group of papers is ”Health”, which includes 8 out of 156 publications. These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease. All the papers working with functional magnetic resonance imaging (fMRI) data were excluded from the analysis, as the dataset was generated as image sequences rather than time series. This key difference in data structure makes them less relevant to the spatio-temporal focus of this review.
@@ 5.4.1 Overview
Applications of GNNs in the ”Health” group have been studied starting from 2021. The reason of this spread may be attributed to two key factors. The first reason is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals which can be used for the timely diagnosis of chronic diseases. The second reason is the spread of the COVID-19 pandemic, which has motivated researchers to investigate and forecast the spread of the virus. Epidemic forecasting is the most common topic among the selected papers. The prediction of epidemics presents two significant challenges. First, their seasonality can vary greatly over time, and they can manifest with different intensities and durations. Second, occasional pandemics can disrupt established patterns even for years. These complexities make the forecasting task more difficult, and machine learning and GNN models have been studied to address this issue. A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to individuals which stands for susceptible (S), infectious (I), and recovered (R). The SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune, and uses differential equations to describe how the disease spreads. In general, some papers use it as benchmark (e.g., Ref. [art_49]).
All the selected papers are published in different journals, but primarily in the field of computer science. The majority of the corresponding authors are affiliated with academic institutions based in China.
@@ 5.4.2 Datasets
Table 13: List of public datasets in the ”Health” group and their corresponding links.
@@
COVID-19 data [art_49] https://www.2019ncov.chinacdc.cn/
COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
@@
Nytimes COVID-19 data [art_50] https://github.com/nytimes/covid-19-data
@@
None of the papers in this section use the same dataset, so it is not possible to compare the results of the proposed models. Most of the datasets are related to the COVID-19 effects and have a daily granularity. In addition, some papers ([art_49], [art_52], [art_92], [art_189]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions).
5.4.3 Proposed models
As for the taxonomy of the proposed models, 5 out of 8 papers use a convolutional GNN architecture. Two papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture. Different approaches are used to define the graph structure. Some papers base their graphs on the geographical distribution of the data (as in Refs. [art_49], [art_52] and [art_92]), and others use models that learn the graph structure directly, as in Refs. [art_48] and [art_50]. There is also a paper that uses a variant of the visibility algorithm [art_14]. In almost all the cases, the graph is static.
In terms of programming languages, this community uses only Python with PyTorch. As for loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks. Tab. LABEL:t:link_codici_health shows the links to the related source codes.
Table 14: List of source codes of the ”Health” models in the review.
@@ 5.4.4 Benchmark models
Tab. LABEL:t:benchmark_models_health lists the benchmark models in the ”Health” group that are discussed in at least two papers.
Table 15: List of benchmark models in the ”Health” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52]
Classical machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
Long-short term memory network (LSTM) [art_49], [art_50]
There are no particularly relevant benchmark models, and none fall within the category of GNNs. This is because GNNs have not been widely studied in the context of health.
@@ 5.4.5 Results
In each paper, the authors claim that the proposed GNN-based model outperforms the benchmarks. The most common error metrics are the MAE and the MSE for forecasting problems and accuracy, sensitivity and specificity for classification problems. A comparison of the accuracy of the different models is not possible due to the lack of common benchmark datasets.
@@ 5.5 Mobility
The topic that includes the largest number of papers identified by the query (39 out of 156) is ”Mobility”, which regards the movement of people. It includes urban traffic, air travel, and bicycle demand, among other applications. Given the large number of articles in this group, a more detailed comparison of benchmarks, models, and results can be provided.
@@ 5.5.1 Overview
Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and preventing urban congestion. The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors that can be used for this purpose. However, making accurate traffic forecasts is challenging due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events. Unlike classical statistical and machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain. The increasing number of papers published each year in this group reflects both the growing research interest in this field and the growing confidence in GNN models. The most commonly investigated areas regard urban traffic ([art_93], [art_95], [art_97], [art_98], [art_99], [art_100], [art_103], [art_105], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_123], [art_124], [art_125], [art_126], [3], [art_128], [art_130], [art_131], [art_132], [art_135]), especially urban traffic flow and urban traffic speed. Among the journals where the selected papers were published, Applied Intelligence by Springer is the most common, with 5 out of 39 papers. This is followed by ACM Transactions on Intelligent Systems and Technology and ACM Transactions on Knowledge Discovery from Data, both published by the Association for Computing Machinery, each with 4 out of 39 papers. Most of the corresponding authors are affiliated with Chinese institutions.
@@ 5.5.2 Datasets
Table 16: List of public datasets in the ”Mobility” group and their corresponding links.
@@
City Bike NYC [art_94], [art_114] https://citibikenyc.com/system-data
DiDi Beijing [art_102] https://github.com/didi/TrafficIndex
@@
Huai’an Bus Transaction [art_101] https://www.pandabus.cn/panda/panda-minibus.html
@@
METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128] https://paperswithcode.com/dataset/metr-la
@@
NYC-Taxi [art_96] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [art_95], [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132] https://paperswithcode.com/dataset/pems04
PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130] https://paperswithcode.com/dataset/pems07
PeMS08 (PeMSD8) [art_95], [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122] https://paperswithcode.com/dataset/pems-bay
@@
RCOTP [art_95] https://developer.ibm.com/exchanges/data/all/airline/
@@
TaxiBJ [art_126] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
@@
The most popular datasets are PeMS ones (http://pems.dot.ca.gov/), provided by the California Transportation Agency Performance Measurement System, which include various versions differing in time range, spatial coverage, and the number of sensors, and METR-LA, collected from loop detectors in Los Angeles County highway. The number of sensors is in the hundreds, and so is the number of nodes in the graph. For a more detailed overview of the benchmark datasets in the traffic domain, see Ref. [review_traffico_con_dataset]. Other common public traffic datasets can be found at https://paperswithcode.com/task/traffic-prediction. In terms of external data, the most commonly used are weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network (e.g., points of interest).
@@
The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead. In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 8 out of 39 papers).
5.5.3 Proposed models

### v2

@@ 5.3.4 Benchmark models
Tab. LABEL:t:benchmark_models_finance displays the benchmark models utilized in the “Finance” group by a minimum of two of the selected papers.
Table 15: List of benchmark models in the “Finance” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_68], [art_375]
Mean reversion (MR) [art_63], [art_65], [art_377]
Non-GNN machine learning methods Convolutional neural network (CNN) [art_59], [art_64], [art_374], [art_379]
Dual-stage attention-based recurrent neural network (DARNN) [art_63], [art_65], [art_377]
Feed-forward neural network (FNN) [art_59], [art_68], [art_375], [art_378]
Gated recurrent unit (GRU) [art_60], [art_68], [art_382]
Long-short term memory network (LSTM) [art_59], [art_62], [art_63], [art_64], [art_65], [art_67], [art_68], [art_374], [art_375], [art_377], [art_379], [art_381]
Support vector machine (SVM) [art_64], [art_68], [art_375], [art_378]
GNN methods Graph convolutional network (GCN) [art_59], [art_60], [art_62], [art_64], [art_68], [art_382]
Graph convolutional network with long short term memory network (GCN+LSTM) [art_65], [art_377]
@@
Hierarchical graph attention network (HATS) [HATS] [art_62], [art_63], [art_65], [art_377]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_62], [art_382]
Spatiotemporal hypergraph convolution network (STHGCN) [STHGCN] [art_63], [art_65], [art_377]
It is evident that there is no consensus regarding the most appropriate benchmark model to use, especially of GNN type. The limited number of GNN-based benchmarks can be attributed to the fact that GNN models have not been extensively explored in the financial domain. The most common benchmark is the LSTM network, which is considered an effective approach for processing long term information in time series. The most commonly used GNN model is a vanilla graph convolutional network (GCN), followed by a more complex hierarchical graph attention network (HATS) proposed by Raehyun et al. in 2019 [HATS]. The HATS model was proposed for predicting stock movements using relational data. It aggregates information from different types of relationships among data and incorporates this information into each stock representation. From an architectural point of view, the model starts with employing LSTM and GRU networks as feature extraction modules. Next, a hierarchical GNN-based architecture is designed to capture the significance of neighboring nodes and relation type across different layers. Finally, the node representations are fed into a task-specific module, which can be implemented as a simple linear transformation layer. The source code for the HATS model is available at https://github.com/dmis-lab/hats.
@@ 5.3.5 Results
The accuracy of the proposed classification models is primarily evaluated in terms of accuracy, precision and F1 score. The only two papers with results which can be compared are Refs. [art_63] and [art_65], which study 758 frequently traded stocks collected from the A-share market in China from January 2013 to December 2019 (available at https://finance.sina.com.cn/stock). These two papers compare the results obtained by using 5, 10 and 20 trading days as records. For sake of simplicity, a comparison of the accuracy of the models in the two papers with 20 days only is provided in Tab. LABEL:t:risultati_a-share.
@@
The two best models are the knowledge graph and graph convolution neural network (KG-GCN) [art_63] and the stock price movement prediction via convolutional network with adversarial hypergraph model (MONEY) [art_65], depending on the metrics considered. The KG-GCN model, proposed by Wang et al. in 2023, first constructs a knowledge graph to represent the semantic relationships and quantify correlations between stocks. Then a community detection module based on GCN node embeddings and K-means clustering is used to group similar stocks together. Finally, the historical prices of similar stocks are fed into LSTM and GRU networks to predict stock price trends. The MONEY model, introduced by Sun et al. in 2023, is composed of a convolutional GNN and an hypergraph convolution network to capture complex relations among similar stocks.
@@ 5.4 Health
The fourth thematic group of papers is “Health”, which includes 15 out of 281 publications. These papers focus on critical aspects of health monitoring, disease modeling, and diagnostic tools, with particular emphasis on the spread and diagnosis of disease. All the papers working with functional magnetic resonance imaging (fMRI) data were excluded from the analysis, as the dataset was generated as image sequences rather than time series. This key difference in data structure makes them less relevant to the spatio-temporal focus of this review.
@@ 5.4.1 Overview
Applications of GNNs to the “Health” thematic group have been studied starting from 2021. The reason of this expansion may be attributed to two key factors. The first factor is the growing presence of sensors in hospitals, which enable the collection of multivariate time series signals, which can be used for the timely diagnosis of chronic diseases. The second factor is the spread of the COVID-19 pandemic, which motivated researchers to investigate and to forecast the spread of the virus. Epidemic forecasting is indeed the most common topic among the selected papers. The prediction of epidemics presents two significant challenges. First, epidemics’ seasonality can vary greatly over time, and they can manifest with different intensities and durations. Second, occasional pandemics can disrupt established patterns even for years. These complexities make the forecasting task more difficult, and non-GNN machine learning, and GNN models themselves, have been studied to address this issue. Moreover, the spread of large-scale human diseases is closely related to both the mobility patterns of populations and the demographic characteristics of residents in different regions [art_413]. Consequently, the application of GNN models to epidemic dynamics appears to be a suitable approach for analyzing and addressing such data. A common reference model for epidemic prediction is the SIR model [modello_SIR], an acronym referring to the state of the health of individuals. The three letters of the acronym stand for susceptible (S), infectious (I), and recovered (R). Namely, the SIR model divides the population into individuals who are susceptible to the disease, those who are infected and can spread it, and those who have recovered and are now immune. It uses a system of differential equations to describe how the disease spreads spatially and changes in time. In general, some papers use it as benchmark (e.g., Ref. [art_49, art_413, art_415]).
Other papers in this section focus on different types of time series, such as electroencephalogram (EEG) data for epilepsy detection (Ref. [art_14]) or emotion recognition (Refs. [art_411, art_412]). While emotion recognition may not seem directly related to the “Health” topic, it is included here because it relies on EEG data. The use of EEG in emotion recognition helps to identify more accurately true emotional states, as there physiological signals cannot be consciously controlled in the same way as facial expressions, tone of voice, or body posture [art_411].
The selected papers were each published in different journals, with one paper per journal, primarily in the field of computer science. The majority of the corresponding authors are affiliated with academic institutions based in China.
@@ 5.4.2 Datasets
Table 17: List of public datasets in the “Health” group and their corresponding links.
@@
African COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
@@
CDC US-State [art_414] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html
Chickenpox [art_410] https://www.kaggle.com/datasets/die9origephit/chickenpox-cases-hungary
EEG DEAP dataset [art_412] http://www.eecs.qmul.ac.uk/mmv/datasets/deap/download.html
@@
European COVID-19 dataset [art_410] https://www.worldometers.info/coronavirus/
ILI IDWR Japan-Prefecture [art_414] https://tinyurl.com/y5dt7stm
ILI US-HHS US-Regional [art_414] https://tinyurl.com/y39tog3h
@@
Nytimes COVID-19 data [art_50], [art_413] https://github.com/nytimes/covid-19-data
OxCGRT dataset [art_416] https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker
@@
Japan COVID-19 [art_415] https://github.com/swsoyee/2019-ncov-japan
@@
SJTU Emotion EEG Dataset (SEED) [art_411], [art_412] https://bcmi.sjtu.edu.cn/home/seed/
@@
Spanish COVID-19 data [art_414] https://dataforgood.fb.com/tools/disease-prevention-maps/
@@
US COVID-19  [^2] This dataset includes multiple sources. [art_415] https://github.com/CSSEGISandData/COVID-19
US COVID-19 [art_416] https://covidtracking.com/data/national
US COVID-19 vaccinations [art_50] https://ourworldindata.org/us-states-vaccinations
@@
Only two pairs of papers in this thematic group use the same dataset, but none of the proposed models yield comparable results. Most of the datasets are related to COVID-19 and have a daily time granularity. In addition, some papers (Refs. [art_49, art_52, art_92, art_189, art_413]) include exogenous variables that can be either context-related (e.g., geography, population, health conditions, vaccinations, economic factors), or specific to a single patient (including personal information such as age, gender, and medical conditions). A paper that undertakes an in-depth study of exogenous variables is Ref. [art_416]. The authors categorize these variables as holiday variables, time categorical variables (e.g., categorical week and month, and the number of days and weeks since the start of the dataset), and policy variables (e.g., school and workplace closing, movement restrictions, face coverings, and vaccination policies). They carry on a significance analysis of these variables, and explore the correlation among the significant ones. The COVID-19 policy dataset (OxCGRT) which includes three years of 23 policy indicators across more than 180 countries is available at https://www.bsg.ox.ac.uk/research/covid-19-government-response-tracker.
5.4.3 Types of models
As for the taxonomy of the proposed models, 11 out of 15 papers use a convolutional GNN architecture. Three papers use a pure attentional model, while another paper presents a hybrid convolutional-attentional architecture. Different approaches are used to define the graph structure. Some papers base their graphs on the geographical distribution of the data (as in Refs. [art_49, art_52, art_92]), and others use models that learn the graph structure directly, as in Refs. [art_48, art_50, art_415, art_416]. There is also a paper that uses a variant of the visibility algorithm [art_14]. In half of the cases, the graph is static, while the dynamic graph approach appears more frequently in the papers from 2024.
In terms of programming languages, the community of the “Health” thematic group uses only Python with PyTorch. As for the coice of the loss functions, MSE is typically used for prediction tasks, and cross entropy is used for classification tasks. Tab. LABEL:t:link_codici_health shows the links to the related source codes.
Table 18: List of source codes of the “Health” models in the review.
@@ 5.4.3 Types of models
Backbone-based Dynamic Spatio-Temporal Graph Neural Network (BDSTGNN) [art_415] https://github.com/JkMaoya/BDSTGNN
@@
Ordinary Differential Equation in Graph Neural Networks (GNNODENet) [art_413] https://github.com/xiongzhangxyq/GNNODENet
Recurrent Graph Gate Fusion Transformers (ReGraFT) [art_416] https://github.com/mfriendly/ReGraFT
@@ 5.4.4 Benchmark models
Tab. LABEL:t:benchmark_models_health lists the benchmark models in the “Health” group that are discussed in at least two papers.
Table 19: List of benchmark models in the “Health” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_48], [art_49], [art_52], [art_410], [art_414], [art_415]
SIR model [modello_SIR] [art_49], [art_413], [art_415]
Non-GNN machine learning methods Feed-forward neural network (FNN) [art_14], [art_48]
Bi-Hemisphere Domain Adversarial Neural Network (Bi-DANN) [Bi-DANN] [art_411], [art_412]
A Novel Bi-Hemispheric Discrepancy Model for EEG Emotion Recognition (BiHDM) [BiHDM] [art_411], [art_412]
Gated recurrent unit (GRU) [art_410], [art_413], [art_415], [art_416]
Long-short term memory network (LSTM) [art_49], [art_50], [art_410], [art_413]
Recurrent neural network (RNN) [art_189], [art_413], [art_414]
Support vector machine (SVM) [art_189], [art_411]
GNN methods Cross-location attention based graph neural network (ColaGNN) [ColaGNN] [art_414], [art_415], [art_416]
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_49], [art_413], [art_414], [art_416]
Dynamical graph convolutional neural networks (DGCNN) [DGCNN] [art_411], [art_412]
Graph attention network (GAT) [GAT] [art_48], [art_413]
Graph convolutional network (GCN) [art_49], [art_413]
Regularized Graph Neural Networks (RGNN) [RGNN] [art_411], [art_412]
Spatio-temporal attention network for pandemic prediction (STAN) [STAN] [art_413], [art_415]
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_413], [art_415]
There are no particularly outstanding benchmark models, and the ones categorized as GNNs appear only in papers from 2024. This is due to the fact that in the past years GNNs have not been extensively explored in the context of health. The most widely used GNN-based benchmark is the diffusion convolutional recurrent neural network (DCRNN), proposed by Li et al. in 2017 for traffic forecasting. The model aims to capture spatial dependencies using bidirectional random walks on the graph (whose adjacency matrix is based on node proximity), and temporal dependencies using an encoder-decoder architecture with scheduled sampling. The source code for the DCRNN model is available at https://github.com/liyaguang/DCRNN.
@@ 5.4.5 Results
In each paper, the authors claim that the proposed GNN-based model outperforms the benchmarks. The most common error metrics are the MAE and the MSE for forecasting problems and accuracy, sensitivity and specificity for classification problems. The forecasting horizon spans from 1 day to more than 10 weeks ahead. A comparison of the accuracy of the different models is not possible due to the lack of common benchmark datasets.
@@ 5.5 Mobility
The thematic group that includes the largest number of papers identified by our query (66 out of 281) is “Mobility”, which regards the movement of people. It includes urban traffic, air travel, and bicycle demand, among other applications. Given the large number of articles in this group, a more detailed comparison of benchmarks, models, and results can be provided.
@@ 5.5.1 Overview
Early traffic prediction is essential for improving the efficiency of transportation systems, helping drivers plan their trips more effectively, and, in the context of cities, preventing urban congestion. The advent of smart cities infrastructures and transportation systems has contributed to the collection of rich data from road sensors. Making accurate traffic forecasts is challenging, due to constantly changing traffic patterns over time and space, as well as the influence of external factors such as weather conditions and special events. Unlike statistical and non-GNN machine learning models, GNNs can be particularly effective for traffic forecasting due to their ability to model complex relationships in spatial and temporal domain. The increasing number of papers published each year in this thematic group reflects both the growing research interest in this field and the growing confidence in GNN models. The most commonly investigated areas regard urban traffic (approximately 70% of the papers), especially urban traffic flow and urban traffic speed. Among the journals where the selected papers were published the most common are Applied Intelligence by Springer, ACM Transactions on Knowledge Discovery from Data, both by the Association for Computing Machinery, and IEEE Transactions on Intelligent Transportation Systems by the Institute of Electrical and Electronics Engineers, with 6 out of 66 papers each. Most of the corresponding authors are affiliated with Chinese institutions.
@@ 5.5.2 Datasets
Table 20: List of public datasets in the “Mobility” group and their corresponding links.
@@
Capital Bike Share [art_446] https://ride.capitalbikeshare.com/system-data
@@
DiDi Beijing [art_102], [art_447] https://github.com/didi/TrafficIndex
@@
Divvy Bike Chicago [art_446] https://ride.divvybikes.com/system-data
Huai’an Bus Transaction [art_452] https://www.pandabus.cn/panda/panda-minibus.html
@@
Lyft bike dataset [art_446] https://www.lyft.com/bikes/bay-wheels/system-data
METR-LA [art_93], [art_99], [art_100], [art_109], [art_110], [art_113], [art_116], [art_120], [art_122], [art_126], [art_128], [art_408], [art_427], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/metr-la
@@
NYC-Bike [art_94], [art_96], [art_114], [art_446], [art_449], [art_451] https://citibikenyc.com/system-data
NYC SafeGraph [art_449] https://www.safegraph.com/academics
NYC-Taxi [art_96], [art_449] https://www1.nyc.gov/site/tlc/about/tlc-trip-record-data.page
PeMS03 [art_108], [art_113], [art_115], [art_117], [art_125], [art_130], [art_419], [art_435], [art_445], [art_453] https://www.kaggle.com/datasets/elmahy/pems-dataset
PeMS04 (PeMSD4) [art_97], [art_99], [art_104], [art_108], [art_113], [art_115], [art_117], [art_122], [art_125], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_424], [art_433], [art_435], [art_441], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems04
PeMS07 (PeMSD7) [art_108], [art_110], [art_113], [art_117], [art_125], [3], [art_130], [art_406], [art_422], [art_424], [art_427], [art_432], [art_435], [art_438], [art_441], [art_444], [art_445], [art_453] https://paperswithcode.com/dataset/pems07
PeMS08 (PeMSD8) [art_97], [art_99], [art_108], [art_115], [art_117], [art_122], [art_125], [art_130], [art_131], [art_132], [art_406], [art_408], [art_424], [art_432], [art_433], [art_435], [art_441], [art_444], [art_445], [art_448], [art_449], [art_452], [art_453] https://paperswithcode.com/dataset/pems08
PEMS-BAY [art_100], [art_109], [art_110], [art_116], [art_120], [art_122], [art_406], [art_408], [art_419], [art_420], [art_422], [art_430], [art_433], [art_435], [art_441], [art_442], [art_444] https://paperswithcode.com/dataset/pems-bay
@@
Seattle loop [art_441] https://github.com/zhiyongc/Seattle-Loop-Data
@@
TaxiBJ [art_126], [art_451] https://www.microsoft.com/en-us/research/publication/forecasting-citywide-crowd-flows-based-big-data/
@@
Urban-core [art_441] https://topis.seoul.go.kr/
Xiecheng delay dataset [art_443] https://pan.baidu.com/s/1dEPyMGh#list/path\$=\$%2F
@@
The most popular datasets used in this group are those provided by the California Transportation Agency Performance Measurement System (PeMS, available at http://pems.dot.ca.gov/), which include various versions differing in time range, spatial coverage, and number of sensors, and the one called METR-LA, collected from loop detectors along the Los Angeles County highway. The number of sensors is in the hundreds, and such is also the number of nodes in the graph. For a more detailed overview of the benchmark datasets in the traffic domain, see Ref. [review_traffico_con_dataset]. Other common public traffic datasets can be found at https://paperswithcode.com/task/traffic-prediction. As to external data, the most commonly used are data on weather conditions, time information (time of day and day of week) and events (e.g., holidays), as well as information about the road network, like points of interest.
@@
The majority of the papers utilize data with a 5-minute granularity, and provide forecasts for horizons ranging from 5 minutes to 1 hour ahead. In addition to imputation of missing values, in almost half of the cases the pre-processing of the data includes data normalization (min-max or Z-score normalization) and the selection of time windows regarding recent, daily and weekly data (as in 10 out of 66 papers).
5.5.3 Types of models